OpenStudy (anonymous):

Where Pn denotes the n-th prime in ascending order, use induction to prove Pn <= 2 ^ ( 2^ ( n )-1 ). @-------College Al…

OpenStudy (anonymous):

this looks tough

OpenStudy (anonymous):

Pn \[\le 2 ^{2^{n}-1}\]

OpenStudy (anonymous):

base case: P(1) ≤ 2 is the first prime smaller than or equal to 2? yes. induction hypothesis: P(k) ≤ 2^(2^k - 1) now we must prove that P(k+1) ≤ 2^(2^(k+1) - 1)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

we can start with the expression 2^(2^(k+1) - 1) = [2^(2^k - 1)]^2 * 2 = 2a^2, where a = 2^(2^k - 1) we know that a ≥ P(k) hence 2a^2 is most definitely ≥ P(k) but since you've taken a number, squared it, and multiplied it by 2, it's definitely going to be bigger than the next prime, too hence 2a^2 ≥ P(k+1) hence 2^(2^(k+1) - 1) ≥ P(k+1)

OpenStudy (anonymous):

hmm, i think the reasoning in my 6th line is a little iffy.

OpenStudy (anonymous):

all the same, i've got a much clearer idea of how to approach this question now. Thank you! :)